# Survey of Graph Neural Network for Internet of Things and NextG Networks

### v1

@@ II-B Architecture of Graph Neural Networks
A simple GNN consists of the following three main stages as shown in Fig. 5.
  • Graph Representaion. A graph is represented as $G=(V,E)$, where $V$ is the set of nodes (vertices) and $E$ is the set of edges (connections) between nodes. Each node $i\in V$ typically has an associated feature vector $\textbf{x}_{i}$ that encodes information about the node. A GNN aims to learn representations of nodes that capture the graph’s structural and semantic information.
@@
  • Message Passing. The core operation in GNNs is message passing, which enables nodes to exchange information with their neighbors in the graph. The message passing step is typically defined in two parts:
@@
    • Aggregation Function: Each node aggregates information from its neighbors by computing an aggregation function (e.g., sum, mean, max) over the features of its neighbors. Mathematically, this can be represented as:
      $m_{i}^{(l+1)}=A(\textbf{h}_{j}^{(l)},\forall j\in N_{i}),$ (1)
@@
      where, $m_{i}^{(l+1)}$ represents the aggregated message at node $i$ at layer $l+1$, $\textbf{h}_{j}$ represents the hidden state of neighbor node $j$ at layer $l$, $A(\cdot)$ represents the aggregation function, $N_{i}$ represents the set of neighbors of node $i$.
@@
      $\textbf{h}_{i}^{(l+1)}$ $=U^{(l)}(\textbf{h}_{j},m_{i}^{(l+1)}),$ (2)
      $\textbf{h}_{i}^{(l+1)}$ $=\sigma(\textbf{W}_{i}^{(l)}\textbf{h}_{i}^{(l)}+\textbf{W}_{j}^{(l)}\textbf{h}_{j}^{(l)})$ (3)
      where, $\textbf{h}_{i}^{(l+1)}$ and $\textbf{h}_{i}^{(l+1)}$ feature vectors of node $i$ at layers $l$ and $l+1$, respectively, $\textbf{W}_{i}^{(l)}$ and $\textbf{W}_{j}^{(l)}$ are the weight matrices for nodes $i$ and $j$ at layers $l$. GNNs typically have multiple layers (or iterations) of message passing and updating. Each layer refines the node representations by considering information from a broader neighborhood in the graph.
@@
II-C Types of Graph Neural Networks
In this section, we will discuss some of the most popular types of variations of GNN architectures namely Graph Convolutional Networks (GCNs), Graph Attention Networks (GAT), Graph Sample and Aggregated (GraphSAGE), Graph Autoencoders (GAE), Graph Isomorphic Networks (GIN), Graph Recurrent Neural Network (Graph-RNN), Graph Capsule Networks (GCaps) and Spatial-Temporal Graph Networks (STGNs).
@@ A LeakyReLU nonlinearity is applied to the Attention mechanism $a$ and can be expressed as
where a is the weight vector for attention mechanism $a$, $(\cdot)^{T}$ and ||||| | are transpose and concatenation operators, respectively. Applying (9) in (8), we get
@@
Graph Sample and AGgregatE (GraphSAGE). GraphSAGE is a graph representation learning technique that aims to generate embeddings for nodes in a graph. GraphSAGE starts by sampling a fixed-size neighborhood for each node in the graph. This neighborhood typically consists of the node itself and its immediate neighbors. The sampling process can be random or based on some specific strategy. For each node, GraphSAGE aggregates information from its sampled neighborhood based on an aggregator function. After aggregating information from the neighborhood, GraphSAGE generates a new embedding for each node. This embedding is learned through a neural network layer, which can be shared across all nodes or specific to each node. GraphSAGE combines techniques from graph sampling and aggregation to create node embeddings that capture both local and global graph structures. The mathematical modeling for GraphSAGE embedding generation [131] can be expressed as follows. First, each node $i$ aggregates the representations of its neighboring nodes into a single vector $\textbf{h}_{N(i)}^{l}$. Then, GraphSAGE concatenates the node’s current representation $\textbf{h}_{i}^{l}$ with the aggregates neighborhood vector $\textbf{h}_{N(i)}^{l}$ which is fed through a fully connected layer with non-linear activation function $\sigma$ to get $\textbf{h}_{i}^{l+1}$. This can be expressed as
@@
$\rm{Loss}$ $=\frac{1}{\mathcal{N}}\sum_{i=0}^{\mathcal{N}}(||S_{\rm{inp},\textit{i}}-S_{\rm{dec},\textit{i}}||^{2})$ (15)
where $\mathcal{N}$ is the number of elements in inp, $S_{\rm{inp}}$ and $S_{\rm{dec}}$ are the sum of elements in inp and dec, respectively, $||(\cdot)||^{2}$denotes the squared Euclidean distance between the input and reconstructed features. GAEs are widely used for Anomaly Detection [139, 138], Wireless Sensor Network Topology Reconstruction [140], Indoor Localization [141], Intrusion Detection [142] and Task Assignment [143].
GNN Variant ML Library Used Related Literature
GCN Tensorflow [144] [145]
PyTorch [146]
GraphSAGE Tensorflow [147] [131]
PyTorch [148]
GAE Tensorflow [149] [150]
PyTorch [151]
GIN Tensorflow [152] [153]
PyTorch [154]
GCapsNet Tensorflow [155] [156]
PyTorch [157]
STGN Tensorflow [158] [159]
PyTorch [160]
TABLE IV: Example GitHub projects and related literature for different GNN variants.
@@
where $\textbf{h}_{i}^{(l)}$ and $\textbf{h}_{i}^{(l+1)}$ are the representation of node $i$ at layers $l$ amd $l+1$, respectively. $N_{i}$ is the set of neighboring nodes of $i$, $\epsilon$ is a learnable or a fixed scalar parameter and $\rm{MLP}$ is a feed-forward multi-layer perceptron. GIN has shown promising results and has been applied to various tasks related to learning and classification [162, 163], network authentication [164], malicious node detection [165], network traffic prediction [166].
Graph Capsule Networks. Graph Capsule Networks (GCapsNet) are a type of neural network architecture inspired by capsule networks originally introduced for image analysis tasks. Graph Capsule Networks extend this concept to graph-structured data, allowing them to handle relational reasoning and hierarchical representations within graph data. In traditional CNNs, convolutional layers typically extract features from images by applying convolutional filters across spatial dimensions. Capsule networks introduce the idea of capsules, which are groups of neurons that represent instantiation parameters of entities within an image, such as object pose, scale, orientation, etc. These capsules allow for more robust and interpretable representations of objects compared to traditional CNNs. An example of GCapsNet called Graph Capsule Convolution Network (GCNN) is introduced in [167]. For node $i$ with node feature value of $x_{0}$ and a set of neighborhood node feature values $\mathcal{F}_{i}=\{x_{0},x_{1},x_{2},...,x_{j}\}$ including node $i$ the output of graph capsule function is $f:\mathbb{R}^{j}\rightarrow\mathbb{R}$ can be expressed as
@@
STGNs integrate techniques from graph neural networks (GNNs) and temporal modeling to effectively capture spatial relationships and temporal dynamics within the data. They typically consist of multiple layers, including graph convolutional layers, temporal modeling layers, and attention mechanisms, to process and learn from the spatio-temporal graph data. STGNs are commonly used for Multi-object Tracking [172], IoT Intrusion Detection [173, 174] and Cellular Traffic Prediction [175, 176].
To further motivate the readers to use GNN in their research domain, Table IV lists some example GitHub projects as well as the related literature that can be referred to as a first step in understanding the design and implementation of different GNN variants.

### v2

@@ II-A Architecture of Graph Neural Networks
A simple GNN consists of the following three main layers namely the input layer, hidden layers, and output layer as shown in Fig. 4. To enable information exchange and aggregation from neighboring nodes in the graph, a more sophisticated GNN is designed using the concept of message passing. With the help of message passing through the graph, GNNs can leverage the local neighborhood information of all the nodes in the graphs and help in refining and updating the global representation of the network. This allows for GNN to capture both the structural information and the features of all the nodes of the graph. The architecture of GNN designed based on message passing is shown in Fig. 5. The working of message passing-based GNN is as follows.
  • Graph Representation. A graph is represented as $G=(V,E)$, where $V$ is the set of nodes (vertices) and $E$ is the set of edges (connections) between nodes. Each node $i\in V$ typically has an associated feature vector $\textbf{x}_{i}$ that encodes information about the node. A GNN aims to learn representations of nodes that capture the graph’s structural and semantic information. The general principle behind generating a graph dataset (i.e., representation). First, identify the problem of interest, this can be energy optimization, reducing latency, improve throughput, among others. The next step is to understand the network topology and the different entities present. The different entities can vary depending on the application. For example, it can be a transmitter, receiver, relay, access point, base station, moving vehicles, etc., Each of these entities is called a node of the graph (refer Fig. 6(a)). The next step is to understand the connections between entities. These connections are called edges. The edges denote the relationship between the entities (or nodes). For example, if one transmitter may be connected to a receiver. Now there is an edge between the transmitter and receiver. This receiver may in turn have connections to two relay nodes, and so on. Now we have the details of the nodes and edges. The next step is to identify the features of the node and the edges. First, let’s begin with the node features. The node features are nothing but the attributes of the node. These features can include details such as node name, ID, location information, speed of travel, transmission power, maximum transmission power, and so on. Next is the edge feature which are attributes of the connection between two nodes. These attributes include link distance, latency, throughput, link capacity, and so on. Finally, all this information is represented in the form of matrices that can be used as the graph representation for the GNN. Figure 7 shows a table of different GNN structure used in representative works discussed later in Sections III-VI [122, 123, 124, 125].
    There are four types of matrices namely feature matrix, adjacency matrix, degree matrix, and weight matrix. The feature matrix represents the 2-D representation of features of all the nodes in the graph. For example, if a graph has $N$ nodes, and each node has $F$ number of features, then the feature matrix F has a dimension of $N\times F$. There are different types of features such as numeric features, categorical features, and binary features, among others. An example of binary features can be a multi-hop network, where each node will have a value of 0 or 1 based on its feature such as it can either be a transmitter (0) or receiver (1), it may be static (1) or mobile (0) and so on. In this case, as an example, the number of features will be 2 and the feature vector will comprise of either 1 or 0 and will be of length 2. If there are 4 nodes in the network, then the feature matrix F will be of size $4\times 2$ (refer Fig. 6(b)). The relationship between the nodes are represented in the form of edge feature matrix (refer Fig. 6(c)). Examples of edge features include the distance between the two nodes, path loss, etc. The adjacency matrix represents the connection between the nodes in a graph and takes a value of 1 or 0 if there is an edge between a pair of nodes or not. It is worth mentioning here that the adjacency matrix is always a square matrix, meaning the number of rows and columns is the same. If there are $N$ nodes in the network, then the shape of the adjacency matrix will always be $N\times N$ (refer Fig. 6(d)). The adjacency matrix can also be a weighted adjacency matrix which is used to represent graphs in which the edges have weights. These weights can represent various attributes such as the strength of a connection, distance, cost, among others. The degree matrix is a diagonal matrix containing information about the degree of each node. Here, the degree of the node indicates the number of terminating edges for a node. For example, a node may have $C$ connections associated with it (refer Fig. 6(e)). Similar to the adjacency matrix, the degree matrix is also a square matrix of shape $N\times N$. Finally, we have the weight matrix which is the trainable parameters used to transform node features during the learning process.
@@
  • Message Passing. The core operation in GNNs is message passing, which enables nodes to exchange information with their neighbors in the graph and can be bi-directional in nature [126, 127]. The message passing step is typically defined in two parts:
@@
    • Aggregation Function: Aggregation refers to the process of collecting information at the node of interest from all the nodes in the neighborhood where neighborhood refers to the set of nodes that are connected to the node $i$. Each node aggregates information from its neighbors by computing an aggregation function (e.g., sum, mean, max) over the features of its neighbors. Mathematically, this can be represented as:
      $\textbf{m}_{N_{i}}^{(l)}$ $=A^{(l)}(\textbf{h}_{j}^{(l)},\forall j\in N_{i}),$ (1)
      $\textbf{m}_{N_{i}}^{(l)}$ $=\sum_{j\in N_{i}}\textbf{h}_{j}^{(l)}$ (2)
@@
      where, $\textbf{m}_{N_{i}}^{(l)}$ represents the aggregated message at node $i$ at layer $l$, $\textbf{h}_{j}^{(l)}$ represents the hidden state of neighbor node $j$ at layer $l$, $A^{(l)}(\cdot)$ represents the aggregation function, $N_{i}$ represents the set of neighbors of node $i$ at layer $l$.
@@
      $\textbf{h}_{i}^{(l+1)}$ $=U^{(l)}(\textbf{h}_{i}^{l},\textbf{m}_{N_{i}}^{(l)}),$ (3)
      $\textbf{h}_{i}^{(l+1)}$ $=\sigma(\textbf{W}_{i}^{(l+1)}\textbf{h}_{i}^{(l)}+\textbf{W}_{j}^{(l+1)}\textbf{m}_{N_{i}}^{(l)}+\mathbf{b}^{(l+1)})$ (4)
      where, $\textbf{h}_{i}^{(l)}$ and $\textbf{h}_{i}^{(l+1)}$ are feature vectors of node $i$ at layers $l$ and $l+1$, respectively, $\textbf{h}_{j}^{(l)}$ is the feature vector of node $j$ at layer $l$ , $\textbf{W}_{i}^{(l+1)}$ and $\textbf{W}_{j}^{(l+1)}$ are the trainable weight matrices for nodes $i$ and $j$ at layers $l+1$, $\textbf{m}_{N_{i}}^{(l)}$ represents the aggregated message at node $i$ at layer $l$, and $\mathbf{b}^{(l+1)}$ is the bias term which is usually a constant added to the product of features and weights and $\sigma$ is the element-wise non-linearity activation function. Activation Functions allow the defining of a nonlinear family of functions to capture the relationship between the input graph data and its representation. Some examples of non-linear activation functions include Sigmoid or Logistic, Hyperbolic Tangent, Rectified Linear Unit (ReLU), LeakyReLU, Parametric ReLU, Exponential Linear Units, Softmax, Scaled Exponential Linear Unit (SELU) and Gaussian Error Linear Unit (GELU), among others. GNNs typically have multiple layers (or iterations) of message passing and updating. Each layer refines the node representations by considering information from a broader neighborhood in the graph.
@@
II-B Types of Graph Neural Networks
In this section, we will discuss some of the most popular types of variations of GNN architectures namely Graph Convolutional Networks (GCNs), Graph Attention Networks (GAT), Graph Sample and Aggregated (GraphSAGE), Graph Autoencoders (GAE), Graph Isomorphic Networks (GIN), Graph Capsule Networks (GCaps), Spatial-Temporal Graph Networks (STGNs) and Graph Reinforcement Learning (GRL). Figure 8 shows the representative architecture of each of the GNN variants to help understand the basic operations.
@@ A LeakyReLU nonlinearity is applied to the Attention mechanism $a$ and can be expressed as
where a is the weight vector for attention mechanism $a$, $(\cdot)^{T}$ and $||$ are transpose and concatenation operators, respectively. Applying (10) in (9), we get
@@
Graph Sample and AGgregatE (GraphSAGE). GraphSAGE is a graph representation learning technique that aims to generate embeddings for nodes in a graph. Embedding refers to the process of representing discrete variables as continuous, low-dimensional vectors. There are 3 types of embedding namely Node Embedding, Edge Embedding, and Graph Embedding. Node embedding involves learning a vector representation for each node in the graph such that nodes with similar network neighborhoods or connectivity patterns have similar embeddings. These embeddings capture the structural and relational information of nodes in the graph. GraphSAGE starts by sampling a fixed-size neighborhood for each node in the graph. This neighborhood typically consists of the node itself and its immediate neighbors. The sampling process can be random or based on some specific strategy. For each node, GraphSAGE aggregates information from its sampled neighborhood based on an aggregator function. After aggregating information from the neighborhood, GraphSAGE generates a new embedding for each node. This embedding is learned through a neural network layer, which can be shared across all nodes or specific to each node. GraphSAGE combines techniques from graph sampling and aggregation to create node embeddings that capture both local and global graph structures. The mathematical modeling for GraphSAGE embedding generation [143] can be expressed as follows. First, each node $i$ aggregates the representations of its neighboring nodes into a single vector $\textbf{h}_{N(i)}^{l}$. Then, GraphSAGE concatenates the node’s current representation $\textbf{h}_{i}^{l}$ with the aggregates neighborhood vector $\textbf{h}_{N(i)}^{l}$ which is fed through a fully connected layer with non-linear activation function $\sigma$ to get $\textbf{h}_{i}^{l+1}$. This can be expressed as
@@
[…]
where $\mathcal{N}$ is the number of elements in inp, $S_{\rm{inp}}$ and $S_{\rm{dec}}$ are the sum of elements in inp and dec, respectively, $(\cdot)||^{2}$denotes the squared Euclidean distance between the input and reconstructed features. GAEs are widely used for Anomaly Detection [151, 150], Wireless Sensor Network Topology Reconstruction [152], Indoor Localization [153], Intrusion Detection [154] and Task Assignment [155].
TABLE IV: Comparison of key features, strengths, and weaknesses of different GNN variants.
GNN Variant Key Features Strengths Weaknesses Dynamic Environment Suitability Expressive Power
Graph Convolutional Network (GCN) Spectral-based method; Aggregates features from neighboring nodes. Simple and effective; Performs well on homophilic graphs. Assumes homogeneity in neighborhoods; Computationally expensive for large graphs. Limited Low
Graph Attention Network (GAT) Employs attention mechanism to weight neighbors; Learnable importance for neighbors.
Handles varying neighborhood sizes.
Differentiates between neighboring nodes; Flexible across diverse graph structures. High computational cost due to attention calculations; Prone to overfitting on small datasets. Limited Low
Graph Sample and AGgregatE (GraphSAGE) Inductive learning method; Samples neighbors for aggregation; Designed for large-scale graphs. Scales well to large and dynamic graphs; Suitable for real-time and inductive settings. Sampling strategy affects performance; May be less accurate than transductive approaches. Suitable Moderate
Graph Autoencoder (GAE) Encoder-decoder architecture; Learns node embeddings unsupervised; Common for link prediction and clustering. Effective for unlabelled graphs; Simple design with strong applications in unsupervised tasks. Depends on well-designed loss functions; Limited in handling complex graph dynamics; May require large datasets. Not ideal Moderate
Graph Isomorphism Network (GIN) Uses MLP for neighborhood aggregation; High expressive power; Designed to distinguish graph structures. Powerful for graph classification; Strong discriminative capability. Requires careful hyperparameter tuning; Computationally intensive; May be over-parameterized for simple graphs. Limited High
Graph Capsule Network (GCapsNet) Uses capsule networks for graphs; Captures hierarchical and part-whole relationships. Handles complex and hierarchical graph structures; Potentially more robust to noise. Complex architecture; High computational cost. Limited High
Spatial-Temporal Graph Network (STGN) Integrates spatial and temporal dependencies; Combines well with RNNs/LSTMs; Designed for dynamic graphs. Effective for time-series prediction on graphs; Models evolving graph behavior well. Requires extensive preprocessing; Training is complex and data-intensive. Best suited High
Graph Reinforcement Learning (GRL) Learns graph policies via reinforcement learning; Can be integrated with various GNNs; Focused on task-specific embeddings. Enables multi-task learning; Facilitates transfer learning; Effective for high-dimensional state spaces. Embedding quality depends heavily on training; Requires large datasets; Complex to tune and deploy. Varies based on underlying GNN architecture Varies based on underlying GNN architecture
@@
where $\textbf{h}_{i}^{(l)}$ and $\textbf{h}_{i}^{(l+1)}$ are the representation of node $i$ at layers $l$ and $l+1$, respectively. $N_{i}$ is the set of neighboring nodes of $i$, $\epsilon$ is a learnable or a fixed scalar parameter and $\rm{MLP}$ is a feed-forward multi-layer perceptron. GIN has shown promising results and has been applied to various tasks related to learning and classification [158, 159], network authentication [160], malicious node detection [161], network traffic prediction [162].
Graph Capsule Networks. Graph Capsule Networks (GCapsNet) are a type of neural network architecture inspired by capsule networks originally introduced for image analysis tasks. Graph Capsule Networks extend this concept to graph-structured data, allowing them to handle relational reasoning and hierarchical representations within graph data. In traditional CNNs, convolutional layers typically extract features from images by applying convolutional filters across spatial dimensions. Capsule networks introduce the idea of capsules, which are groups of neurons that represent instantiation parameters of entities within an image, such as object pose, scale, orientation, etc. These capsules allow for more robust and interpretable representations of objects compared to traditional CNNs. An example of GCapsNet called Graph Capsule Convolution Neural Network (GCNN) is introduced in [163]. For node $i$ with node feature value of $x_{0}$ and a set of neighborhood node feature values $\mathcal{F}_{i}=\{x_{0},x_{1},x_{2},...,x_{j}\}$ including node $i$ the output of graph capsule function is $f:\mathbb{R}^{j}\rightarrow\mathbb{R}$ can be expressed as
@@
STGNs integrate techniques from graph neural networks (GNNs) and temporal modeling to effectively capture spatial relationships and temporal dynamics within the data. They typically consist of multiple layers, including graph convolutional layers, temporal modeling layers, and attention mechanisms, to process and learn from the spatio-temporal graph data. A single convolutional layer can be expressed as in Eq. 7, which can then be modified to account for spatial and temporal dependencies as follows [168].
$\textbf{h}^{(l+1)}=\sigma(\textbf{D}_{t}^{-\frac{1}{2}}\textbf{A}_{t}\textbf{D}_{t}^{-\frac{1}{2}}\textbf{h}_{s}^{(l)}\textbf{W}_{t}^{(l)})$ (20)
$\textbf{h}_{s}^{(l)}=\textbf{D}_{s}^{-\frac{1}{2}}\textbf{A}_{s}\textbf{D}_{s}^{-\frac{1}{2}}\textbf{h}^{(l)}\textbf{W}_{s}^{(l)}$ (21)
where $\textbf{h}^{(l)}$, $\textbf{h}^{(l+1)}$ are the feature vectors, at layers $l$ and $l+1$, respectively, $\textbf{h}_{s}^{(l)}$ is the spatial feature vector, $\textbf{A}_{t}$ and $\textbf{A}_{s}$ are temporal and spatial adjacency matrices, $\textbf{D}_{t}$ is the degree matrix of $\textbf{A}_{t}$, and $\textbf{D}_{s}$ is the degree matrix of $\textbf{A}_{s}$. $\textbf{W}_{t}^{(l)}$ and $\textbf{W}_{s}^{(l)}$ are temporal and spatial weight matrix of layer $l$, and $\sigma$ is an activation function, such as ReLU or sigmoid. Finally, STGN (ie.e, STGCN) can be expressed as
$\textbf{h}_{t}^{(l+1)}=\sigma(\textbf{D}_{s}^{-\frac{1}{2}}\textbf{A}_{s}\textbf{D}_{s}^{-\frac{1}{2}}\textbf{h}^{(l)}\textbf{W}_{s}^{(l)}\textbf{W}_{t}^{(l)})$ (22)
STGNs are commonly used for Multi-object Tracking [169], IoT Intrusion Detection [170, 171] and Cellular Traffic Prediction [172, 173].
Graph Reinforcement Learning. Graph Reinforcement Learning (GRL) is a special type of GNN that uses RL for real-time interaction with the environment and decision-making. RL is used with GNN due to its ability to learn by interacting with the environment making it suitable for complex, dynamic, and uncertain environments where there is no prior knowledge of the network to train the GNN model. GNNs are commonly combined with RL techniques such as Q-Learning [174], REINFORCE [175], Actor-Critic [176], among others. This has motivated researchers to apply GRL for various problems such as Multiaccess Edge Computing [177], Joint Cruise Control and Task Offloading [112], Resource Allocation [131, 132]. In addition to the combination of simple GNN with RL, there is also the combination of different variants of GNNs with RL such as GCN-RL [178], GAT-RL [139] and GraphSAGE-RL [179].
In Table IV, we compare the key features, strengths, weaknesses, suitability for dynamic environment and expressive power of different GNN variants to help readers choose the best variant based on the application’s needs.